The length of a rectangular garden is 9 feet longer than its width. If the garden's perimeter is 206 feet, what is the  of the garden in square feet? 

Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 206ft
L + w = 103ft
(w+9) + w = 103ft
2w = 94ft
w = 47ft and L = 56ft
A = Lw = (47ft)(56ft)= 2,632 ft^2